solve for x:(log3 X)^2 + log3 X^2 +1=0
Answer by stanbon(57967)

:(log3 X)^2 + log3 X^2 +1=0
(log3 x )^2 + 2log3 x + 1 = 0
Factor:
(log3 x +1)^2 = 0
log3 x = -1
x^-1 = 3
x = 1/3 (with multiplicity two)
====================================
Cheers,
Stan H.